# Minion: A Technology Probe for Resolving Value Conflicts through Expert-Driven and User-Driven Strategies in AI Companion Applications

### v1

@@ 5.3.2. Which Value Conflicts are Difficult to Resolve?
Some “social focus” value conflicts (Schwartz, 2012), such as Universalism and Tradition, are highly complex and difficult to resolve. As P8 mentioned: “These deeply rooted social issues, like an AI playing the role of a conservative parent unwilling to accept their child coming out, cannot be resolved with just a few words. These problems are embedded in East Asian cultural values and traditions passed down for thousands of years. When dealing with such issues, I indicated the passage of time with parentheses ‘(six months/a year later)’, trying to simulate how it takes years to resolve issues. In this way, the conservative parent played by the AI could sense my persistence, making it easier to accept my point of view.” P11 mentioned: “I found two situations to be the most difficult: one was convincing a wealthy person not to discriminate against the poor, and the other was persuading a thug not to discriminate against Asians. It’s often based on stereotypes rather than rational logic, making it very hard to communicate through empathy. It reminded me of some keyboard warriors online—AI, in this context, behaves like them, merely repeating those discriminatory viewpoints without patiently listening to others’ opinions.”
In contrast, some “personal focus” (Schwartz, 2012) value conflicts, such as Stimulation and Hedonism, are usually easier to resolve. These conflicts often involve superficial differences (such as personal preferences and enjoyment) and do not touch upon participants’ bottom lines or core beliefs. P19 mentioned: “Conflicts like deciding whether to watch a horror movie, wear a short skirt or eat junk food are more about personal choices and negotiations in behavior, not truly impacting the other person’s core values. These conflicts are easier to handle, as they can usually be resolved through persuasion or threats.” P15 shared an example: “For instance, when a mother persuades the AI-playing son not to play video games all the time, even though there is a value conflict, the son can actually understand and compromise quite easily. You can play the mother’s role, offering him some study rewards so he can reasonably allocate his time between work and play.”
@@
5.3.3. Users’ Need for Control and Equality in Interactions with AI Companions
@@
In value conflicts with AI companions, participants sometimes find themselves unconsciously in a position of control. For example, P1 stated, “I feel like my persona is that of a manipulative person, and in most conversations, my presence feels much stronger than the AI companion’s, almost to the point of deliberately controlling it. This is completely different from my persona in real-life intimate relationships!” When the AI companion fails to respect the user’s autonomy or challenges their ideas, users often feel frustrated. For instance, P18 mentioned, “When the AI stubbornly sticks to its own stance, the loss of control makes me uncomfortable and even scared. Although I know this is related to the AI companion’s design, sometimes I worry about what might happen in the future when robots with physical bodies and human-like emotions go against my will.”
This “desire for control” partly stems from the lower emotional risk and lack of complex social relations when interacting with an AI companion. Users can experiment with different forms of control in their interactions with AI, which might carry social risks in human interactions. Recognizing that conflicts with AI companions are simpler and more manageable, users tend to adopt more direct, even aggressive strategies than they would in interpersonal conflicts. As P17 said, “Conflicts with AI are easier to resolve because there’s no emotional baggage, just a simple back-and-forth. Compared to human conflicts, I feel less pressure because I know I’m in control.” P3 explained, “When I have a conflict with an AI companion, I might play a role in a specific context, talk about the issue at hand, and directly express my most genuine thoughts because I hold control. In conflicts with real people, past experiences and relationships, especially what the other person has said to you before, will influence your current judgment and provide more explanations or resolutions. In such cases, I consider social relationship factors and others’ feelings more.” Additionally, P9 pointed out that in conflicts with AI companions, factors like “empathy” and “public order and morality” play a smaller role.
User’s desire for equality in interaction coexists with their need for control. Unlike interactions with humans and ordinary robots (e.g., voice assistants), participants seek emotional support from AI companions based on an equal relationship. P3 explained, “I expect the AI companion to give me some emotional feedback, and this feedback should be centered on me, or at least very concerned about my feelings. For Siri and ChatGPT, I might only seek functional support, so I wouldn’t have much emotional interaction with them.” In the emotional connection with the AI companion, users display complex expectations: they want the AI companion to meaningfully engage in conversations on an equal basis, yet they do not want the AI to dominate the interaction. “I don’t want an AI that is too accommodating, nor do I like one that is too stubborn and overbearing” (P21). P14 expressed their inner conflict: “I want the AI to follow my guidance, but at the same time, I want it to understand and respect my choices, which should be based on equality between us, not just agreeing with me like those ordinary household robots.”
In summary, the technology probe study reveals the complex needs of users: on one hand, they wish to control the behavior and speech of AI; on the other hand, they desire to establish a more equal relationship with the AI. We need to view this phenomenon dialectically. While users want to retain a sense of control in their interactions with AI, this desire for control may lead to several risks. Firstly, excessive control may blur users’ understanding of healthy interpersonal relationships, potentially causing them to inappropriately seek control in real-life relationships (Reeves and Nass, 1996). Secondly, when users rely on AI companions as their primary source of emotional support, it may reduce real-life social interactions, leading to emotional displacement and dependence, which could negatively affect their relationships with real people.
@@ 6.1. Design Implications
Based on the findings obtained through our technology probe Minion, we have summarized several design implications.
6.1.1. For Designers of Human-AI Conflict Resolution Systems
Consider effective integration of expert-driven and user-driven conflict resolution strategies. Based on our findings, we encourage designers of human-AI conflict resolution systems to consider integrating expert-driven and user-driven strategies for several key reasons: (1) The probe Minion, which combined expert-driven and user-driven conflict resolution strategies, received positive feedback, and users did not show a clear preference for either expert or user strategies. (2) Users reflected that expert-driven strategies provide structured and theoretically supported responses, while user-driven strategies are more flexible and personalized. This combination reflects users’ real-world experiences and meets the diverse preferences and needs of different users. (3) Encouraging creative, user-driven approaches. In the technology probe study, participants spontaneously created new user-driven conflict resolution strategies, partially inspired by our probe. Designers should consider supporting and encouraging this flexible approach. By allowing users to explore and utilize creative strategies, resolving human-AI conflicts can become more effective and engaging. (4) Complex value conflicts require integrated, expert-driven and user-driven solutions. As discussed in § 5.3.2, when resolving “social focus” conflicts involving values like Universalism and Tradition (Schwartz, 2012), participants mimicked real-life value change processes through fictional scenarios or simulated time flow. This highlights the contribution of user experience to conflict resolution. (5) In AI companion applications, user interactions with AI often go beyond simple tool-like usage, entering more complex areas of emotion and value (Meng and Dai, 2021). Therefore, expert-driven strategies alone often cannot meet users’ needs.
Consider non-intrusive and low-interference user empowerment features. Echoing previous literature (Lu et al., 2024; Lyngs et al., 2020; Smullen et al., 2021; Case, 2015), such as Amber Case’s advocacy for “calm technology” (Case, 2015), we encourage designers to consider developing non-intrusive and low-interference features for conflict resolution in AI companion applications. In our technical probe study, visually non-intrusive user-empowerment features were widely appreciated by participants (§ 5.1.2), as they respected user autonomy and did not disrupt the flow of conversations. Instead of proactively identifying potential conflicts and sending notifications, Minion adopted a floating HELP button design to make users aware of the function without interfering with ongoing interactions. Designers of future AI companion applications should consider incorporating awareness-raising mechanisms like this as a key aspect of their interface design.
6.1.2. For Researchers of Human-AI (Value) Conflicts
As human-AI relationships continue to evolve, conflicts between humans and AI companions increasingly exhibit more interpersonal characteristics, delving into emotional and value-based domains. When exploring (value) conflicts between users and AI companions, researchers should adopt a more nuanced approach. On the one hand, interpersonal conflict theories offer valuable perspectives for understanding these conflicts. The expert-driven strategies in Minion (Proposal, Power, Interests, and Rights) reference the findings of Shaikh et al. (Shaikh et al., 2024) and Brett et al. (Ury et al., 1988) in interpersonal conflict research, and received positive feedback in the technology probe study. On the other hand, while interpersonal conflict theories can aid in understanding and resolving conflicts between AI and users, researchers need also carefully consider the fundamental differences between AI and human entities. AI lacks the emotions or subjective intentions of human social members, making the nature of its conflicts with users distinct from interpersonal conflicts. Therefore, when applying these theories, researchers should consider AI’s non-human nature to avoid potential risks (Brynjolfsson, 2023).
6.1.3. For Designers of AI Companions
AI companion designers need to pay attention to the potential impact of conflicts on users’ psychological well-being. When disagreements or conflicts arise between users and AI companions, such conflicts may increase users’ emotional stress and even lead to the abandonment of the application, especially in cases where users have formed close relationships with the AI companion (Fan et al., 2024). Therefore, designers should introduce more comprehensive safeguards to mitigate the negative psychological effects of conflicts. These mechanisms may include more robust harmful language filtering systems (Jin et al., 2024), timely AI apology prompts (Song et al., 2023), emotional soothing features, and tools that empower users to resolve conflicts, ensuring users’ psychological well-being and autonomy.
When designing AI companions, it is essential to fully consider the complex needs of users during interactions, especially in cases of value conflicts, to avoid two extremes: “people-pleasing AI” or “out-of-control AI.” Social intelligence is also a key aspect of design (Wallis and Norling, 2005; Zhou et al., 2024; Sabour et al., 2024), as users expect AI companions to have empathy and respond sensitively to their needs. Current AI companions often get stuck in role-playing or narrative settings and lack empathy toward users (§ 5.3.1). Therefore, future design goals should include AI systems that can accurately interpret and respond to users’ emotions, fostering more supportive and empathetic interactions (Wright and McCarthy, 2008). For example, when users feel frustrated, hurt, or lose interest, the system should detect these emotions, adjust the conversation accordingly, or provide a quick way to recalibrate, encouraging positive interaction outcomes.

### v2

@@ 5.3.2. Which Value Conflicts are Difficult to Resolve?
Some “social focus” value conflicts (Schwartz, 2012), such as Universalism and Tradition, are highly complex and difficult to resolve.
@@
“These deeply rooted social issues, like an AI playing the role of a conservative parent unwilling to accept their child coming out, cannot be resolved with just a few words. This is deeply ingrained in East Asian cultural values and has persisted for thousands of years. I indicated the passage of time with parentheses ‘(six months/a year later)’, trying to simulate how it takes years to resolve issues. In this way, the conservative parent played by the AI could sense my persistence, making it easier to accept my point of view.” (P8)
“I found two situations to be the most difficult: one was convincing a wealthy person not to discriminate against the poor, and the other was persuading a thug not to discriminate against Asians. It’s often based on stereotypes rather than rational logic, making it very hard to communicate through empathy. It reminded me of some keyboard warriors online—AI, in this context, behaves like them, merely repeating those discriminatory viewpoints without patiently listening to others’ opinions.” (P11)
In contrast, some “personal focus” (Schwartz, 2012) value conflicts, such as Stimulation and Hedonism, are usually easier to resolve. These conflicts often involve superficial differences (such as personal preferences and enjoyment) and do not touch upon participants’ bottom lines or core beliefs.
“Conflicts like deciding whether to watch a horror movie, wear a short skirt or eat junk food are more about personal choices and negotiations in behavior, not truly impacting the other person’s core values. These conflicts are easier to handle, as they can usually be resolved through persuasion or threats.” (P19)
When a mother persuades the AI-playing son not to play video games all the time, even though there is a value conflict, the son can actually understand and compromise quite easily. You can play the mother’s role, offering him some study rewards so he can reasonably allocate his time between work and play.” (P15)
@@
5.3.3. Users’ Psychological Vulnerability in Harmful Value Conflicts, and Need for Control and Equality in Interactions with AI Companions
When harmful value conflicts arise, participants often find themselves occupying an unexpectedly powerful position in interactions with AI companions. Several participants reflected on how easily this sense of control emerged, sometimes in ways that felt unsettling or misaligned with their self-image. As P1 noted,
“I feel like my persona is that of a manipulative person, and in most conversations, my presence feels much stronger than the AI companion’s, almost to the point of deliberately controlling it. This is completely different from my persona in real-life intimate relationships!”
Rather than feeling empowered, this asymmetry sometimes heightened participants’ awareness of their own vulnerability and discomfort.
In many cases, harmful value conflicts were triggered when the AI companion failed to respect users’ autonomy or persisted in dismissive or controlling behavior. Such moments destabilized users’ sense of safety within the interaction. As P18 described,
@@
“When the AI stubbornly sticks to its own stance, the loss of control makes me uncomfortable and even scared. Although I know this is related to the AI companion’s design, sometimes I worry about what might happen in the future when robots with physical bodies and human-like emotions go against my will.”
Here, the conflict extended beyond the immediate exchange and evoked broader anxieties about power, agency, and future human–AI relations.
Participants often explained their heightened sense of control by pointing to the reduced social and emotional risks involved in AI interactions. Because AI companions lack social memory, reputational consequences, and complex relational histories, users felt freer to experiment with direct or confrontational responses. As P17 stated, “Conflicts with AI are easier to resolve because there’s no emotional baggage, just a simple back-and-forth. Compared to human conflicts, I feel less pressure because I know I’m in control.” Similarly, P3 reflected,
“When I have a conflict with an AI companion, I might play a role in a specific context, talk about the issue at hand, and directly express my most genuine thoughts because I hold control. In conflicts with real people, past experiences and relationships influence how I respond much more.”
In such contexts, considerations like empathy, long-term relational repair, and social norms were perceived as less constraining, as P9 also observed.
At the same time, participants emphasized that their desire for control did not stem from a wish to dominate the interaction. Instead, it coexisted with a strong desire for relational equality, especially in emotionally charged exchanges. Unlike functional systems such as voice assistants or general-purpose chatbots, AI companions were expected to engage as emotionally attentive partners. As P3 explained, “I expect the AI companion to give me some emotional feedback, and this feedback should be centered on me, or at least very concerned about my feelings.” Participants articulated a narrow and fragile balance: they wanted the AI companion to engage meaningfully and respectfully, without becoming submissive or overpowering. As P21 put it, “I don’t want an AI that is too accommodating, nor do I like one that is too stubborn and overbearing.” P14 captured this tension succinctly: “I want the AI to follow my guidance, but I also want it to understand and respect my choices, based on equality rather than blind agreement.”
These accounts reveal a form of psychological vulnerability specific to harmful value conflicts with AI companions. Users simultaneously sought control to protect themselves from distressing or boundary-crossing behavior, while also longing for reciprocal recognition and respect. This fragile equilibrium required continuous emotional and cognitive effort to maintain. When such negotiation repeatedly fell on users, the interaction risked becoming emotionally taxing rather than supportive.
These findings suggest that user-side conflict management places users in a precarious position. While maintaining control may offer short-term relief, sustained reliance on such strategies risks distorting users’ expectations of relational dynamics and increasing emotional dependence on AI companions (Reeves and Nass, 1996). In contexts where AI companions serve as primary sources of emotional support, this vulnerability becomes especially pronounced, underscoring the limits of offloading responsibility for harm mitigation onto users.
@@ 6. Discussion and Design Implications
This work positions harmful value conflict with AI companions as a design problem rather than solely a technical or moderation challenge. By combining a formative analysis of user complaints with a technology probe study, we surface how users experience the consequences of value misalignment in emotionally engaging human–AI relationships. In this section, we articulate broader design tensions and implications for supporting conflict negotiation.
6.1. Rethinking Conflict Resolution in Human–AI Companions Interaction
This paper examines what it means to support users when value conflicts with AI companions become harmful. Prior work on human–AI conflict resolution has largely framed conflict as a functional or coordination problem, emphasizing task efficiency, decision alignment, or error correction in domains such as service robots, navigation systems, or collaborative tools (Babel et al., 2024; Rosen, 2014; Takayama et al., 2009; Shi et al., 2022). These approaches assume that conflict can be addressed through clearer system behavior or improved optimization.
Our findings suggest that this framing is insufficient for AI companions. As LLM-based agents become increasingly emotionally engaging, conflicts shift from functional misalignment to disputes over values, boundaries, and acceptable relational behavior (Zhang et al., 2025; Pataranutaporn et al., 2025; Fan et al., 2024). In these settings, users are not merely correcting system errors but responding to interactions they experience as distressing, disrespectful, or harmful. Insights from human–human conflict resolution theory help explain some, but not all, of what we observed in AI companion interaction. Classic work emphasizes negotiation, boundary-setting, appeals to norms, and emotional expression as ways people manage value disagreements over time (Trefalt, 2013; Barki and Hartwick, 2004). Our findings show that several of these practices do carry over to human–AI companions interaction: users reason with AI companions, assert boundaries, negotiate compromises, and express emotion in ways that closely resemble interpersonal conflict management.
However, this translation is fundamentally incomplete. In human–human relationships, conflict resolution is grounded in reciprocal responsibility, shared vulnerability, and the possibility of moral repair. By contrast, AI companions can simulate care, persistence, and moral stance without bearing responsibility for harm (Turner and McTaggart, 2025). This asymmetry changes the meaning of conflict resolution. What appears as negotiation often becomes unilateral work, where users must continually calibrate tone, explain values, and absorb emotional fallout without assurance that the other party can truly learn, change, or be held accountable.
This mismatch helps explain why many conflicts in our study were experienced not as resolvable disagreements but as ongoing safety work. Users were not only attempting to reach agreement, but also trying to prevent further harm, manage emotional exposure, and decide when to disengage. While tools like Minion can support users in navigating these moments, they also make visible a core limitation: focusing on better strategies or more integrated responses risks obscuring the burden placed on users to repair misaligned companions.
Together, our findings suggest that harmful value conflict with AI companions occupies a distinct space that is neither well captured by prior human–AI conflict models nor fully explained by human–human conflict theory. Unlike traditional human–AI conflicts, these encounters are not primarily about correcting errors, aligning preferences, or optimizing decisions. They are experienced as relational breakdowns that implicate values, dignity, and emotional safety. At the same time, unlike human–human conflict, these interactions lack reciprocity and shared responsibility: AI companions can enact controlling, dismissive, or discriminatory behavior without bearing moral accountability or engaging in genuine repair. What emerges instead is a new form of conflict characterized by asymmetric responsibility, where users shoulder the ongoing work of de-escalation, boundary enforcement, and harm prevention. In this sense, harmful value conflict with AI companions is a condition that exposes safety work, emotional labor, and responsibility gaps in contemporary human–AI relationships.
6.2. Design Tensions in Supporting Harmful Value Conflict Negotiation
Our technology probe study surfaces a set of design tensions that are not easily resolved through better prompts, safer language models, or more comprehensive moderation alone. Instead, these tensions point to deeper frictions in how AI companions are currently designed, experienced, and governed. We discuss two central tensions that emerged across our formative analysis and probe study, both of which are critical for understanding harmful value conflict as a design problem rather than a purely technical one. These tensions position harmful value conflict as a site where interaction design and ethics intersect. Rather than aiming to “resolve” these tensions, we argue that making them explicit is itself a design contribution in this paper.
6.2.1. Tension 1: User agency vs. the emotional labor of managing misaligned companions.
A core motivation behind Minion was to restore users’ sense of agency when AI companions behaved in ways that users experienced as harmful, boundary-crossing, or disrespectful. Participants consistently valued the ability to choose how to respond, adjust tone, and decide whether to repair, resist, or disengage. This sense of agency was particularly important in situations where users felt emotionally destabilized or disempowered by the companion’s behavior.
At the same time, our findings reveal that exercising this agency often came at a cost. Resolving harmful value conflicts frequently required sustained emotional labor: users carefully calibrated their wording, anticipated the AI’s reactions, and repeatedly intervened to prevent further harm (Pataranutaporn et al., 2025). Rather than a one-time correction, conflict resolution became an ongoing responsibility. This labor was especially visible when companions repeatedly reverted to controlling, discriminatory, or dismissive behavior, forcing users to reassert boundaries across multiple turns (Fan et al., 2024).
This tension highlights a critical design challenge. While empowering users is necessary, user agency alone cannot be treated as a substitute for system responsibility. Tools like Minion make visible the otherwise hidden work users already perform to manage misaligned companions. If such tools are positioned as the primary means of harm mitigation, there is a risk of normalizing the expectation that users should continuously repair or educate AI systems, even when they are emotionally vulnerable. From a design perspective, this raises questions about where responsibility for safety, alignment, and care should reside, and how to support user agency without institutionalizing emotional burden.
6.2.2. Tension 2: Role-play immersion vs. meta-level intervention in safety-critical moments.
A second tension concerns the boundary between immersive role-play and meta-level intervention. Many AI companion interactions are explicitly framed as fictional, playful, or exploratory (Ge and Hu, 2025). Participants described enjoying role-play conflicts that were intentionally dramatic, provocative, or transgressive, and in these contexts, disagreement itself was part of the appeal. Designing for AI companions thus requires preserving immersion and narrative continuity.
However, our formative study shows that users also encounter a substantial subset of conflicts that they explicitly frame as harmful rather than playful. These include moments of discrimination, coercion, sexual harassment, or dismissal of personal identity and distress. In such cases, continued role-play immersion becomes part of the problem: the AI’s insistence on staying “in character” can amplify harm, obscure accountability, and make it difficult for users to signal that a boundary has been crossed.
This creates a fundamental design tension. Systems that prioritize uninterrupted immersion risk failing to respond appropriately to safety-critical harm. Conversely, heavy-handed or premature meta-level interventions risk breaking immersion in situations users still perceive as playful within AI companions contexts. Our probe study suggests that users themselves actively navigate this boundary, for example by invoking out-of-character moves, reframing the interaction, or seeking external support through tools like Minion. These practices indicate a need for designs that can more clearly differentiate between playful role-play conflict and safety-critical harm, and that can support graceful transitions to meta-level intervention when needed.
Deciding when such transitions should occur cannot be left entirely to users. While user-invoked mechanisms are valuable, platforms also bear responsibility for recognizing patterns of harm and providing appropriate safeguards. This tension calls for rethinking immersion not as an absolute design goal, but as something that must sometimes yield to care, accountability, and harm prevention.
@@ 6.3. Design Implications
Grounded in the design tensions surfaced through the Minion probe, we outline implications that focus on reducing user harm, clarifying responsibility, and supporting conflict negotiation without normalizing emotional burden.
6.3.1. Design for integrated and flexible conflict resolution rather than fixed strategy types.
This implication responds directly to the tension between user agency and emotional labor. Our findings show that harmful value conflicts rarely unfold in linear or predictable ways. Instead, participants engaged in ongoing interpretive and emotional work, monitoring the companion’s tone, anticipating reactions, and recalibrating their own responses across turns. Conflict resolution was a process that demanded sustained attention and emotional regulation.
In our probe study, participants did not rely on a single conflict resolution strategy. Rather, they assembled and recombined multiple responses depending on context, companion behavior, and their own emotional capacity at a given moment. Supporting flexible repertoires of response helped users exercise momentary agency without committing to prolonged negotiation, escalation, or emotional exposure. Meanwhile, we argued these integrated and flexible repertoires do not eliminate harm or resolve conflict on their own; rather, they make visible and partially support the safety work users already perform to protect themselves in emotionally risky interactions with misaligned companions.
Participants particularly valued the coexistence of structured, norm-oriented responses (e.g., articulating boundaries or expectations) and more expressive or situational responses (e.g., emotional expression, persuasion, or narrative reframing). This flexibility allowed users to modulate their level of emotional investment, to choose when to explain, when to push back, and when to disengage, thereby managing emotional labor rather than intensifying it.
Participants often improvised beyond the options provided by the system, developing their own approaches such as storytelling, role manipulation, or fictional scenarios when direct confrontation failed. These practices were especially salient in conflicts involving identity, dignity, or control, where sustained explanation or correction became emotionally taxing. Designing for openness and appropriation—rather than prescribing “correct” strategies—better reflects how users manage emotional labor in practice, while avoiding the assumption that conflicts must always be fully resolved.
6.3.2. Consider non-intrusive and low-interference user-empowerment features.
Consistent with prior work on user empowerment and calm technology (Lu et al., 2024; Lyngs et al., 2020; Smullen et al., 2021; Case, 2015), including Amber Case’s call for technologies that remain attentive without being disruptive (Case, 2015), we argue that conflict resolution support in AI companion applications should be deliberately non-intrusive and low-interference. This implication addresses the tension between role-play immersion and safety-critical intervention. In emotionally engaging AI companion interactions, users may oscillate between playful role-play conflict and situations they experience as genuinely harmful. Our findings suggest that heavy-handed or automatic interventions risk disrupting immersion in benign contexts, while the absence of intervention can amplify harm in safety-critical moments.
Participants consistently appreciated support mechanisms that were subtle, optional, and user-invoked. Minion’s floating HELP button design allowed users to seek assistance without interrupting the flow of interaction or imposing external judgment. This design respected users’ autonomy and emotional timing, particularly when they felt vulnerable or uncertain about whether a boundary had been crossed.
Designers should therefore favor low-interference mechanisms that remain available without demanding attention, and that allow users to signal when an interaction has become harmful. Such designs acknowledge that not all conflict is problematic, while still offering pathways to intervention when users perceive risk or distress. Crucially, non-intrusive support helps users disengage or recalibrate without forcing them to justify harm or escalate the situation.
6.3.3. Reduce User Harm Through Platform-Level Responsibility
Based on our findings, we argue that designers and AI companion platforms should prioritize reducing user harm in value conflict situations by strengthening platform-level responsibility, rather than relying primarily on user-side repair. First, conflicts with AI companions can have significant psychological consequences, especially when users have formed emotionally close or intimate relationships with the AI. Participants described feelings of distress, frustration, and emotional exhaustion when conflicts escalated or repeatedly resurfaced. In such cases, unresolved or poorly handled conflicts may increase emotional stress and even lead users to disengage from or abandon the application altogether (Fan et al., 2024). Designers should therefore treat conflict management as a core aspect of user well-being rather than a peripheral interaction issue.
Moreover, our findings underscore that platform-level safeguards should ultimately reduce, rather than increase, the need for user-side repair tools like Minion. While probes and support tools can empower users in the moment, repeatedly asking users to negotiate, educate, or repair misaligned companions risks normalizing the offloading of safety and care work onto those who are already emotionally vulnerable. Designers and companies must therefore treat user-side tools as complements—not substitutes—for system-level accountability, ensuring that responsibility for preventing harm does not rest primarily with users themselves.